# Altermagnetic superconducting diode effect

### v1

@@ I Introduction
The broad class of magnetically ordered states referred to as “altermagnets” (AMs) has recently become a subject of notable interest and has motivated a significant amount of theoretical Yuan et al. (2020); Šmejkal et al. (2022a, b); Ahn et al. (2019); Bhowal and Spaldin (2022); Cuono et al. (2023); Guo et al. (2023); Maier and Okamoto (2023); Mazin (2023); Oganesyan et al. (2001); Sato et al. (2023); Steward et al. (2023); Turek (2022); Ouassou et al. (2023); Mazin et al. (2021); Hayami et al. (2019); Šmejkal et al. (2020); Brekke et al. (2023); Das et al. (2023); Fakhredine et al. (2023); Fang et al. (2023); Leeb et al. (2023); Mazin et al. (2023); Šmejkal et al. (2023); Sun and Linder (2023); Gao et al. (2023); Beenakker and Vakhtel (2023); Zhou et al. (2024); Fernandes et al. (2023); Sun et al. (2023); Zhang et al. (2023a); Brekke et al. (2023); Giil and Linder (2023); Zhu et al. (2023); Li and Liu (2023); Ghorashi et al. (2023); Cheng and Sun (2024); Antonenko et al. (2024); Yu et al. (2024); Wei et al. (2023); McClarty and Rau (2023) and experimental Aoyama and Ohgushi (2023); Bai et al. (2023); Krempaský et al. (2023); Lee et al. (2023); Reimers et al. (2023); Feng et al. (2022) studies, probing into its nature and emergence. Just as with any form of magnetic order, time-reversal symmetry is broken in an AM. However, as opposed to ferromagnets, AMs do not exhibit a finite net magnetic moment, which is guaranteed by a magnetic point symmetry. This means that the state is invariant under the product $\Theta g_{m}$ of time-reversal $\Theta$ and a unitary lattice symmetry $g_{m}$. Unlike a collinear antiferromagnet, $g_{m}$ is not lattice translation for an AM, where the reversal of all local magnet moments resulting from application of $\Theta$ cannot be undone by a lattice translation $T_{\boldsymbol{R}}$. Instead, $g_{m}$ is a point group symmetry such as a rotation or reflection symmetry of the underlying crystalline lattice.
@@
Concurrently, non-reciprocal transport in superconductors and superconducting junctions, particularly the superconducting diode effect (SDE), has garnered various experimental Jiang et al. (1994); Ando et al. (2020); Lyu et al. (2021); Du et al. (2023); Sundaresh et al. (2023); Kealhofer et al. (2023); Hou et al. (2023); Chen et al. (2023); Gupta et al. (2022); Baumgartner et al. (2022); Banerjee et al. (2023); Pal et al. (2022); Kim et al. (2023); Gupta et al. (2022); Turini et al. (2022); Bauriedl et al. (2022); Wu et al. (2022); Diez-Merida et al. (2021); Golod and Krasnov (2022); Chiles et al. (2022); Zhang et al. (2023b); Shin et al. (2021); Jeon et al. (2022); Kealhofer et al. (2023); Hou et al. (2023); Lin et al. (2022); Anwar et al. (2022); Narita et al. (2022); Gutfreund et al. (2023); Zhao et al. (2023) and theoretical Daido and Yanase (2022); Daido et al. (2022); Yuan and Fu (2022); He et al. (2022a); Ilić and Bergeret (2022); Scammell et al. (2022); Zinkl et al. (2022); Zhang et al. (2022); He et al. (2022b); Zhai et al. (2022); Jiang et al. (2022); Kokkeler et al. (2022); Debnath and Dutta (2024); Chazono et al. (2022); Vodolazov and Peeters (2005); de Picoli et al. (2023); Kochan et al. (2023); Ikeda et al. (2022); Tanaka et al. (2022); Wang et al. (2022); Haenel and Can (2022); Legg et al. (2023); Cuozzo et al. (2023); Souto et al. (2022); Cheng and Sun (2023); Steiner et al. (2022); Costa et al. (2023); Wei et al. (2022); Legg et al. (2022); Karabassov et al. (2022); Hu et al. (2023); Wu et al. (2023); Nunchot and Yanase (2024); Daido and Yanase (2023); Souto et al. (2023) attention. The SDE refers to a difference in the critical current $J_{c}(\hat{\boldsymbol{n}})$ in opposite directions, $\hat{\boldsymbol{n}}$ and $-\hat{\boldsymbol{n}}$. In that case, applying a current $J$ with $\min(J_{c}(\pm\hat{\boldsymbol{n}}))<J<\max(J_{c}(\pm\hat{\boldsymbol{n}}))$, leads to superconducting transport along one and resistive behavior along the opposite direction, yielding a superconducting analogue of a semi-conducting diode. On top of fundamental theoretical questions, the rich space of potential applicability of such non-reciprocal responses on modern-day electronics and devices renders the SDE an exciting topic to delve into from an experimental and technological standpoint.
@@ II.1 Diode effect efficiency
To set the stage for our symmetry discussion, we begin by first outlining the basic procedure for the calculation of the critical current. Denoting the superconducting order parameter for Cooper pairs with center-of-mass momentum $\boldsymbol{q}$ by $\Delta_{\boldsymbol{q}}$, we expand the change of the free-energy as a result of superconductivity $\delta\mathcal{F}_{\text{S}}:=\mathcal{F}[\Delta_{\boldsymbol{q}}]-\mathcal{F}[0]$ up to quadratic order in $\Delta_{\boldsymbol{q}}$,
$\delta\mathcal{F}_{\text{S}}\sim\sum_{\boldsymbol{q}}a^{\text{S}}_{\boldsymbol{q}}\,|\Delta_{\boldsymbol{q}}|^{2}+b^{\text{S}}\sum_{\boldsymbol{q}_{i}}\Delta^{*}_{\boldsymbol{q}_{1}}\Delta^{*}_{\boldsymbol{q}_{2}}\Delta_{\boldsymbol{q}_{3}}\Delta_{\boldsymbol{q}_{4}}\delta_{\boldsymbol{q}_{1}+\boldsymbol{q}_{2},\boldsymbol{q}_{3}+\boldsymbol{q}_{4}},$ (1)
@@
Clearly, Eq. (3) is obeyed for all $\hat{\boldsymbol{n}}$ if $g$ is inversion ($I$) or two-fold rotation along the out-of-plane ($\hat{\boldsymbol{z}}$) direction $C_{2z}$ for a 2D system; these symmetries are thus required to be broken to allow for a finite SDE. While the following constraints are readily generalized to three spatial dimensions, we next focus for notational simplicity on a 2D system in the $xy$ plane. If $g$ is a two-fold rotational symmetry, $C_{2\hat{\boldsymbol{m}}}$, along an in-plane direction $\hat{\boldsymbol{m}}$ or a mirror-plane symmetry, $\sigma_{\hat{\boldsymbol{m}}}$, spanned by $\hat{\boldsymbol{z}}$ and $\hat{\boldsymbol{m}}$, Eq. (3) implies that $\eta(\hat{\boldsymbol{n}})=0$ for $\hat{\boldsymbol{n}}$ perpendicular to $\hat{\boldsymbol{m}}$. In other words, the zeros of $\eta(\hat{\boldsymbol{n}})$ are pinned along high-symmetry directions if the point group contains two-fold rotations along in-plane directions or mirror planes perpendicular to the system plane. Note that rotations with degrees higher than two do not pin the zeros.
@@
Equation (4) contains the special case of preserved time-reversal symmetry, $g=\mathbbm{1}$ for which $g\hat{\boldsymbol{n}}=\hat{\boldsymbol{n}}$ always holds, implying $\eta(\hat{\boldsymbol{n}})=0$ for all $\hat{\boldsymbol{n}}$; this just recovers the often-stated fact that $\Theta$ needs to be broken to get a SDE. However, Eq. (4) also shows that zeros of the angle-resolved SDE efficiency are pinned along high-symmetry directions if the system exhibits a magnetic reflection symmetry, $\Theta\sigma_{\hat{\boldsymbol{m}}}$, or magnetic rotational symmetry, $\Theta C_{2\hat{\boldsymbol{m}}}$; in both cases, it holds $\eta(\hat{\boldsymbol{m}})=0$. We will see examples of all of these constraints in the explicit model calculations of Sec. III below.
@@ II.2 AM-induced SDE
$\mathcal{H}_{0}=\sum_{\boldsymbol{k}}c^{\dagger}_{\boldsymbol{k},s}\left[s_{0}\epsilon_{\boldsymbol{k}}+\boldsymbol{g}_{\boldsymbol{k}}\cdot\boldsymbol{s}+\boldsymbol{N}_{\boldsymbol{k}}\cdot\boldsymbol{s}\right]_{s,s^{\prime}}c_{\boldsymbol{k},s^{\prime}},$ (5)
@@
The different possible physical realizations of this class of models are illustrated in Fig. 1. Either the AM term $\boldsymbol{N}_{\boldsymbol{k}}$ will be proximity-induced in superconducting systems, see Fig. 1(a), or vice versa—a proximitized superconductor induces pairing in a metallic AM material [Fig. 1(b)]. Finally, our analysis also applies if both AM and superconductivity arise spontaneously in the same electron liquid, see (c). Although not essential for the realization of the following effects, we here assumed that the proximitizing materials are placed symmetrically on both sides such that an electric field, $E_{z}$, can be used to further reduce the symmetries in situ. Our symmetry analysis and minimal-model study here still applies when the proximitizing material is only present on one side. In that case, the terms induced by $E_{z}\neq 0$ in the models below will already be present for $E_{z}=0$.
@@
$\boldsymbol{S}=\sum_{\boldsymbol{k}}c^{\dagger}_{\boldsymbol{k},s}\boldsymbol{s}_{s,s^{\prime}}c_{\boldsymbol{k},s^{\prime}},$ (7)
has to vanish. For instance, an AM order parameter transforming under $A_{g}$ of $\mathcal{G}_{0}=D_{2h}$ will lead to a residual $C_{2z}$ rotation symmetry, implying that $\braket{S_{x}}=\braket{S_{y}}=0$. The additional $\sigma_{v}$ mirror symmetry further implies that also $\braket{S_{z}}$ vanishes. In the first three columns of Table 1, we list all such combinations of $\mathcal{G}_{0}$ and AM order parameters, some of which were already tabulated in Ref. Fernandes et al., 2023. To organize the list, the first 11 lines refer to the AM order parameters for centrosymmetric point groups $\mathcal{G}_{0}$; interestingly, we find that, in all cases, these AM order parameters will continue to be proper AMs with $\braket{\boldsymbol{S}}=0$ when an electric field is applied, $E_{z}\neq 0$, which reduces the point groups to $\mathcal{G}_{E_{z}}\leq\mathcal{G}_{0}$. Combing back to our previous example, $\mathcal{G}_{0}=D_{2h}$ will be reduced to $\mathcal{G}_{E_{z}}=C_{2v}$. It still contains the two symmetries used above to show that there is no finite magnetization. The remaining lines of Table 1 refer to the AM order parameters in the 2D non-centrosymmetric point groups that are not reached by applying $E_{z}$ in centrosymmetric point groups. In all those cases, reducing the point group further by application of $E_{z}$ leads to a finite magnetization.
@@
Table 1: Summary of possible AM order parameters $\boldsymbol{N}_{\boldsymbol{k}}$ in 2D models of the form of Eq. (5). For each point group $\mathcal{G}_{0}$, we list all the irreducible representations (IR) that can give rise to an AM; the respective form of $\boldsymbol{N}_{\boldsymbol{k}}$ for small $\boldsymbol{k}$ is shown in the third column, where $\tau$ and $R_{\varphi}$ are, respectively, a real-valued constant and a rotation matrix along the third direction not fixed by symmetry. Applying an electric field $E_{z}$ reduces the point group to $\mathcal{G}_{E_{z}}$. In case of the centrosymmetric $\mathcal{G}_{0}$ (first 11 lines), all order parameters remain AMs (as indicated in the fourth column), characterized by zero net magnetization; this is not the case for the non-centrosymmetric $\mathcal{G}_{0}$ (line 12-20). In the column SDE, we indicate whether the respective AM will give rise to an SDE (in the centrosymmetryic point groups, only possible for $E_{z}\neq 0$). We also indicate whether the angle-resolved SDE efficiency $\eta(\hat{\boldsymbol{n}})$ has zeros along generic directions $\hat{\boldsymbol{n}}_{0}$ (and their $C_{3z}$-related partners) or along crystalline axes, $\hat{\boldsymbol{e}}_{x,y}$. For $D_{3}$ and $D_{6}$, we introduced $\hat{\boldsymbol{n}}^{i}_{E_{z}}$ with $\hat{\boldsymbol{n}}^{i}_{E_{z}=0}=\hat{\boldsymbol{e}}_{i}$, $i=x,y$, and pointing along a generic direction for $E_{z}\neq 0$. The last column shows candidate materials, based on their conduction type—metallic (M) or insulating (I)—focusing on those cases where an SDE is present.

### v2

@@ I Introduction
The broad class of magnetically ordered states referred to as “altermagnets” (AMs) has recently become a subject of notable interest and has motivated a significant amount of theoretical Yuan et al. (2020); Šmejkal et al. (2022a, b); Ahn et al. (2019); Bhowal and Spaldin (2022); Cuono et al. (2023); Guo et al. (2023); Maier and Okamoto (2023); Mazin (2023); Oganesyan et al. (2001); Sato et al. (2023); Steward et al. (2023); Turek (2022); Ouassou et al. (2023); Mazin et al. (2021); Hayami et al. (2019); Šmejkal et al. (2020); Brekke et al. (2023); Das et al. (2023); Fakhredine et al. (2023); Fang et al. (2023); Leeb et al. (2023); Mazin et al. (2023); Šmejkal et al. (2023); Sun and Linder (2023); Gao et al. (2023); Beenakker and Vakhtel (2023); Zhou et al. (2024); Fernandes et al. (2023); Sun et al. (2023); Zhang et al. (2023a); Brekke et al. (2023); Giil and Linder (2023); Zhu et al. (2023); Li and Liu (2023); Ghorashi et al. (2023); Cheng and Sun (2024); Antonenko et al. (2024); Yu et al. (2024); Wei et al. (2023); McClarty and Rau (2023) and experimental Aoyama and Ohgushi (2023); Bai et al. (2023); Krempaský et al. (2023); Lee et al. (2023); Reimers et al. (2023); Feng et al. (2022); Bose et al. (2022) studies, probing into its nature and emergence. Just as with any form of magnetic order, time-reversal symmetry is broken in an AM. However, as opposed to ferromagnets, AMs do not exhibit a finite net magnetic moment, which is guaranteed by a magnetic point symmetry. This means that the state is invariant under the product $\Theta g_{m}$ of time-reversal $\Theta$ and a unitary lattice symmetry $g_{m}$. Unlike a collinear antiferromagnet, $g_{m}$ is not lattice translation for an AM, where the reversal of all local magnet moments resulting from application of $\Theta$ cannot be undone by a lattice translation $T_{\boldsymbol{R}}$. Instead, $g_{m}$ is a point group symmetry such as a rotation or reflection symmetry of the underlying crystalline lattice.
@@
Concurrently, non-reciprocal transport in superconductors and superconducting junctions, particularly the superconducting diode effect (SDE), has garnered various experimental Jiang et al. (1994); Ando et al. (2020); Lyu et al. (2021); Du et al. (2023); Sundaresh et al. (2023); Kealhofer et al. (2023); Hou et al. (2023); Chen et al. (2023); Gupta et al. (2022); Baumgartner et al. (2022); Banerjee et al. (2023); Pal et al. (2022); Kim et al. (2023); Gupta et al. (2022); Turini et al. (2022); Bauriedl et al. (2022); Wu et al. (2022); Diez-Merida et al. (2021); Golod and Krasnov (2022); Chiles et al. (2022); Zhang et al. (2023b); Shin et al. (2021); Jeon et al. (2022); Kealhofer et al. (2023); Hou et al. (2023); Lin et al. (2022); Anwar et al. (2022); Narita et al. (2022); Gutfreund et al. (2023); Zhao et al. (2023) and theoretical Daido and Yanase (2022); Daido et al. (2022); Yuan and Fu (2022); He et al. (2022a); Ilić and Bergeret (2022); Scammell et al. (2022); Zinkl et al. (2022); Zhang et al. (2022); He et al. (2022b); Zhai et al. (2022); Jiang et al. (2022); Kokkeler et al. (2022); Debnath and Dutta (2024); Chazono et al. (2022); Vodolazov and Peeters (2005); de Picoli et al. (2023); Kochan et al. (2023); Ikeda et al. (2022); Tanaka et al. (2022); Wang et al. (2022); Haenel and Can (2022); Legg et al. (2023); Cuozzo et al. (2023); Souto et al. (2022); Cheng and Sun (2023); Steiner et al. (2022); Costa et al. (2023); Wei et al. (2022); Legg et al. (2022); Karabassov et al. (2022); Hu et al. (2023); Wu et al. (2023); Nunchot and Yanase (2024); Daido and Yanase (2023); Souto et al. (2023); Cayao et al. (2024) attention. The SDE refers to a difference in the critical current $J_{c}(\hat{\boldsymbol{n}})$ in opposite directions, $\hat{\boldsymbol{n}}$ and $-\hat{\boldsymbol{n}}$. In that case, applying a current $J$ with $\min(J_{c}(\pm\hat{\boldsymbol{n}}))<J<\max(J_{c}(\pm\hat{\boldsymbol{n}}))$, leads to superconducting transport along one and resistive behavior along the opposite direction, yielding a superconducting analogue of a semi-conducting diode. On top of fundamental theoretical questions, the rich space of potential applicability of such non-reciprocal responses on modern-day electronics and devices renders the SDE an exciting topic to delve into from an experimental and technological standpoint.
@@ II.1 Diode effect efficiency
To set the stage for our symmetry discussion, we begin by first outlining the basic procedure for the calculation of the critical current. Denoting the superconducting order parameter for Cooper pairs with center-of-mass momentum $\boldsymbol{q}$ by $\Delta_{\boldsymbol{q}}$, we expand the change of the free-energy as a result of superconductivity $\delta\mathcal{F}_{\text{S}}:=\mathcal{F}[\Delta_{\boldsymbol{q}}]-\mathcal{F}[0]$ up to quartic order in $\Delta_{\boldsymbol{q}}$,
$\delta\mathcal{F}_{\text{S}}\sim\sum_{\boldsymbol{q}}a^{\text{S}}_{\boldsymbol{q}}\,|\Delta_{\boldsymbol{q}}|^{2}+b^{\text{S}}\sum_{\boldsymbol{q}_{i}}\Delta^{*}_{\boldsymbol{q}_{1}}\Delta^{*}_{\boldsymbol{q}_{2}}\Delta^{\phantom{*}}_{\boldsymbol{q}_{3}}\Delta^{\phantom{*}}_{\boldsymbol{q}_{4}}\delta_{\boldsymbol{q}_{1}+\boldsymbol{q}_{2},\boldsymbol{q}_{3}+\boldsymbol{q}_{4}},$ (1)
@@
Clearly, Eq. (3) is obeyed for all $\hat{\boldsymbol{n}}$ if $g$ is inversion ($I$) or two-fold rotation along the out-of-plane ($\hat{\boldsymbol{z}}$) direction $C_{2z}$ for a 2D system; these symmetries are thus required to be broken to allow for a finite SDE. While the following constraints are readily generalized to three spatial dimensions, we first focus for notational simplicity on a 2D system in the $xy$ plane. If $g$ is a two-fold rotational symmetry, $C_{2\hat{\boldsymbol{m}}}$, along an in-plane direction $\hat{\boldsymbol{m}}$ or a mirror-plane symmetry, $\sigma_{\hat{\boldsymbol{m}}}$, spanned by $\hat{\boldsymbol{z}}$ and $\hat{\boldsymbol{m}}$, Eq. (3) implies that $\eta(\hat{\boldsymbol{n}})=0$ for $\hat{\boldsymbol{n}}$ perpendicular to $\hat{\boldsymbol{m}}$. In other words, the zeros of $\eta(\hat{\boldsymbol{n}})$ are pinned along high-symmetry directions if the point group contains two-fold rotations along in-plane directions or mirror planes perpendicular to the system plane. Note that rotations with degrees higher than two do not pin the zeros.
@@
Equation (4) contains the special case of preserved time-reversal symmetry, $g_{m}=\mathbbm{1}$ for which $g_{m}\hat{\boldsymbol{n}}=\hat{\boldsymbol{n}}$ always holds, implying $\eta(\hat{\boldsymbol{n}})=0$ for all $\hat{\boldsymbol{n}}$; this just recovers the often-stated fact that $\Theta$ needs to be broken to get a SDE. However, Eq. (4) also shows that zeros of the angle-resolved SDE efficiency are pinned along high-symmetry directions if the system exhibits a magnetic reflection symmetry, $\Theta\sigma_{\hat{\boldsymbol{m}}}$, or magnetic rotational symmetry, $\Theta C_{2\hat{\boldsymbol{m}}}$; in both cases, it holds $\eta(\hat{\boldsymbol{m}})=0$. We will see examples of all of these constraints in the explicit model calculations of Sec. III below.
We finally point out that Eqs. (3) and (4) also apply for the three-dimensional case. As such, we see that a two-fold non-magnetic rotation (mirror) symmetry lead to line (point) nodes for $\eta(\hat{\boldsymbol{n}})$ on the sphere $\{\hat{\boldsymbol{n}}\in\mathbb{R}^{3}|\hat{\boldsymbol{n}}^{2}=1\}$, while magnetic rotations (mirror) symmetries imply point (line) nodes.
@@ II.2 AM-induced SDE
$\mathcal{H}_{0}=\sum_{\boldsymbol{k}}c^{\dagger}_{\boldsymbol{k},s}\left[s_{0}\epsilon_{\boldsymbol{k}}+\boldsymbol{g}_{\boldsymbol{k}}\cdot\boldsymbol{s}+\boldsymbol{N}_{\boldsymbol{k}}\cdot\boldsymbol{s}\right]_{s,s^{\prime}}c^{\phantom{\dagger}}_{\boldsymbol{k},s^{\prime}},$ (5)
@@
The different possible physical realizations of this class of models are illustrated in Fig. 1. Either the AM term $\boldsymbol{N}_{\boldsymbol{k}}$ will be proximity-induced in superconducting systems, see Fig. 1(a), or vice versa—a proximitized superconductor induces pairing in a metallic AM material [Fig. 1(b)]. Although not essential for the realization of the following effects, we here assumed that the proximitizing materials are placed symmetrically on both sides such that an electric field, $E_{z}$, can be used to further reduce the symmetries in situ. Our symmetry analysis and minimal-model study here still applies when the proximitizing material is only present on one side. In that case, the terms induced by $E_{z}\neq 0$ in the models below will already be present for $E_{z}=0$. Finally, our analysis also applies if both AM and superconductivity arise spontaneously in the same electron liquid, see Fig. 1(c). We note in passing that this is the only scenario which would allow to realize altermagnetic diodes as a bulk effect in three dimensions and where the backaction mechanism of Ref. Banerjee and Scheurer, 2024 can apply, possibly giving rise to enhanced $\eta$.
@@
$\boldsymbol{S}=\sum_{\boldsymbol{k}}c^{\dagger}_{\boldsymbol{k},s}\boldsymbol{s}_{s,s^{\prime}}c^{\phantom{\dagger}}_{\boldsymbol{k},s^{\prime}},$ (7)
has to vanish. For instance, an AM order parameter transforming under $A_{g}$ of $\mathcal{G}_{0}=D_{2h}$ will lead to a residual $C_{2z}$ rotation symmetry, implying that $\braket{S_{x}}=\braket{S_{y}}=0$. The additional $\sigma_{v}$ mirror symmetry further implies that also $\braket{S_{z}}$ vanishes. In the first three columns of Table 1, we list all such combinations of $\mathcal{G}_{0}$ and AM order parameters, some of which were already tabulated in Ref. Fernandes et al., 2023. To organize the list, the first 11 lines refer to the AM order parameters for centrosymmetric point groups $\mathcal{G}_{0}$; interestingly, we find that, in all cases, these AM order parameters will continue to be proper AMs with $\braket{\boldsymbol{S}}=0$ when an electric field is applied, $E_{z}\neq 0$, which reduces the point groups to $\mathcal{G}_{E_{z}}\leq\mathcal{G}_{0}$. Coming back to our previous example, $\mathcal{G}_{0}=D_{2h}$ will be reduced to $\mathcal{G}_{E_{z}}=C_{2v}$. It still contains the two symmetries used above to show that there is no finite magnetization. The remaining lines of Table 1 refer to the AM order parameters in the 2D non-centrosymmetric point groups that are not reached by applying $E_{z}$ in centrosymmetric point groups. In all those cases, reducing the point group further by application of $E_{z}$ leads to a finite magnetization.
@@
Table 1: Summary of possible AM order parameters $\boldsymbol{N}_{\boldsymbol{k}}$ in 2D models of the form of Eq. (5). For each point group $\mathcal{G}_{0}$, we list all the irreducible representations (IR) that can give rise to an AM; the respective form of $\boldsymbol{N}_{\boldsymbol{k}}$ for small $\boldsymbol{k}$ is shown in the third column, where $\tau$ and $R_{\varphi}$ are, respectively, a real-valued constant and a rotation matrix along the third direction not fixed by symmetry. Applying an electric field $E_{z}$ reduces the point group to $\mathcal{G}_{E_{z}}$. In case of the centrosymmetric $\mathcal{G}_{0}$ (first 11 lines), all order parameters remain AMs (as indicated in the fourth column), characterized by zero net magnetization; this is not the case for the non-centrosymmetric $\mathcal{G}_{0}$ (line 12-20). In the column SDE, we indicate whether the respective AM will give rise to an SDE (in the centrosymmetric point groups, only possible for $E_{z}\neq 0$). We also indicate whether the angle-resolved SDE efficiency $\eta(\hat{\boldsymbol{n}})$ has zeros along generic directions $\hat{\boldsymbol{n}}_{0}$ (and their $C_{3z}$-related partners) or along crystalline axes, $\hat{\boldsymbol{e}}_{x,y}$. For $D_{3}$ and $D_{6}$, we introduced $\hat{\boldsymbol{n}}^{i}_{E_{z}}$ with $\hat{\boldsymbol{n}}^{i}_{E_{z}=0}=\hat{\boldsymbol{e}}_{i}$, $i=x,y$, and pointing along a generic direction for $E_{z}\neq 0$. The last column shows candidate materials, based on their conduction type—metallic (M) or insulating (I)—focusing on those cases where an SDE is present.